OpenStudy (buttermequeasy):

Given the function f(x) = 2(x + 10), find x if f(x) = 24. 68 7 17 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

you have but to solve this for your answer: 24=2(x+10)

OpenStudy (anonymous):

24=2(x+10) 12 = x+10 x = 2